The symbols Vf, V.. bols vf, V az oy F, and V x F are defined by Of = grad f 7. F = div F VXF = curl F After eq. (3.12) is me number 4 parts a b and c
Answer #1

Bare D DXF so curl = DXF = 1 i į k xy z -(x+y-3) al-(xty - 2 1 - (xsy-3)) = 2 ( 7 th 2 (3) - 2 lata pody - 4-3 - 1 2 (2)=0 dx